What is the probability of rolling a 4 or an even number on a number cube?
That depends on how many sides the cube has but if it was a 6 sided cube then the probability of getting a four is 1 to 6 and even is 3 to 6.
Since 4 is even, then the result set of 4 intersect with the result set of even, so the probability of rolling a 4 or an even number is 1 in 2, or 0.5.
There are 13 clubs (including the seven of clubs), and 3 other sevens (from the other suits). So that makes 16 cards that are clubs or sevens. Thus, the probability of drawing a club or a seven from a standard 52 card deck is 16/52 = 4/13 (about 31%).
1. He takes 6 steps forward and 5 steps backward
2. He takes 5 steps forward and 6 steps backward
consider 1:
p1 = (11C6)(0.4)^6*(0.6)^5
for 2:
p2 = 11C5(0.4)^5*(0.6)^6
total probability = p1 + p2
What are the deck odds for a pack of playing cards?
There are twenty odd cards in a deck of cards. This is a math problem.
What are the odds of two players being dealt pocket aces in Texas hold 'em?
Roughly speaking, one in 49000 chances.
- Bdx
Correct.
What is probability of getting head or tail when tossing a coin?
The probability of getting a head or a tail is, for all practical purposes, equal to 1. The only other possibility is that the coin stands on its edge - technically possible but so very unlikely that it can safely be ignored.
Percentage between 415 and 515 standard deviation 100 and mean is 515?
34.1% with the normal distribution.
What is the theoretical probability of rolling two dice 100 times?
The theoretical probability is 1. Lots of people have done it in the course of their lives so it is an event that has happened and will happen again.
41 (13 + 13 + 13 + 2)
What percentage chance to you have at being murdered?
The answer depends, among other factors, on what your occupation is and where in the world you live. For most people, the probability is negligible.
What are your chances of conceiving twins after you miscarried twins?
I was told by my doctor, that the chances of having multiples is very likely once you've conceived multiples in the past. And if you have multiples, they will skip a generation and then their kids will likely have multiples. Pretty cool.
Ancient Olympic Games were extended from a one day festival with a few events to a five day event with wrestling, boxing, horse racing, jumping, javelin, and chariot racing.
You want: P(even) * P(even) * P(even) * P(even). P(even) = 3/6 or 1/2. Therefore probability even on 4 rolls is: 1/2 * 1/2 * 1/2 * 1/2 = 1/16 = 0.0625.
The probability is 7/36.
A die is rolled 5 times what is the probability of rolling exactly one 2?
Possible outcomes of rolling a die 5 times = 6^5
Number of outcomes including rolling exactly one 2: 5^5 (5*5^4)
5^5/6^5 = 3125/7776 ~= 0.4019
What does it mean if a number has a probability of 1?
If a number has a probability of 1 it means it is certain to occur.